OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (shamim):

equation of ur line will b\[y-y _{1}=m(x-x _{1})\]

OpenStudy (shamim):

given slope\[m=4\]

OpenStudy (shamim):

\[x _{1}=-3,y _{1}=-1\]

OpenStudy (shamim):

did u get it

OpenStudy (anonymous):

y-4x-11=0 will be ur ans

OpenStudy (shamim):

ya

OpenStudy (shamim):

u can give medal

OpenStudy (shamim):

just click on best answer

OpenStudy (anonymous):

mr shamim collecting medals ..

OpenStudy (shamim):

lol.....

OpenStudy (shamim):

i solved ur math. so i desire medal. is not it

OpenStudy (anonymous):

@@ ss u deserve it ...